# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v1

@@ 3.2 Preliminary Results
Despite its wide use in domain adaptation and instruction tuning, LoRA is not yet explored under the task of LLM unlearning to the best of our knowledge. Therefore, we first share empirical results from low-rank adapting LLMs using gradient difference (Maini et al., 2024; Liu et al., 2022a) as our objective to motivate our approach. For this experiment, we follow previous work (Jang et al., 2023) unlearn 32 text sequences from the Training Data Extraction Challenge from GPT-Neo-1.3B model pretrained on the Pile dataset (Gao et al., 2020). More information on the details on the setup can be found in Section 4 and the Appendix.
@@
Figure 2 shows the results. Notably, vanilla LoRA suffers from lack of plasticity and ends up failing to sufficiently unlearn ${\mathcal{D}}_{f}$ within 20 epochs. When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in LM-eval and Dialogue performances. In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the compute-efficiency of LoRA.
3.3 Inverted Hinge Loss: Novel Loss Function for Unlearning in LLMs
@@
Motivation: Inherent Limitation of GA. We analyze the inherent issues of GA from the perspective of its gradients. The output layer of a language model is a softmax layer that outputs probabilities over the vocabulary. Let ${y}_{t}$ be the logits (pre-softmax activations) produced by the LLM model for the $t$-th token, and let $V$ be the vocabulary size. The probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$ where $y_{t}^{(x_{t})}$is the logit corresponding to the true token $x_{t}$ and $y_{t}^{(v)}$ is the logit corresponding to the $v$-th token in the vocabulary. When we use $\mathcal{L}_{GA}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is:
@@
$\frac{\partial\log(p_{\theta}(x_{t}|x_{<t}))}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
Given this, the derivative of $\mathcal{L}_{GA}({\bm{x}})$ with respect to the logits at each $t$ becomes $\nabla_{y_{t}^{(v)}}\mathcal{L}_{GA}({\bm{x}})={\partial\log(p_{\theta}(x_{t}|x_{<t}))}/{\partial y_{t}^{(v)}}$. From this derivative of GA, we can confirm its unlearning mechanism and inherent issues. First, when $v=x_{t}$, GA decreases the prediction score corresponding to the true token while increasing the prediction score for all other tokens. This demonstrates that GA performs unlearning by encouraging the model to predict tokens different from the true token.
However, despite the successful experimental results of GA (Jang et al., 2023) in LLMs, our gradient analysis reveals its inherent issues. When $v=x_{t}$, it increases the prediction scores of all tokens except the true token. As a result, GA has problems: 1) the loss not converging and becoming unbounded, and 2) unnecessary additional forgetting due to the increase in logits for all other tokens, 3) a large vocabulary size induces gradients to spread out across all other logits, making gradient updates for unlearning inefficient.
@@
Inverted Hinge Loss (IHL). Considering the limitations of GA discussed above, our motivation is to develop a simple yet novel loss function for unlearning in LLMs. The new loss function aims to effectively achieve unlearning by decreasing the prediction score of the true token while increasing the prediction scores of only a minimal number of other tokens. Inspired by the hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as shown below:
@@
Considering the probability $p_{\theta}(x_{t}|x_{<t})$ given by the softmax function, the derivative of $\mathcal{L}_{\text{IH}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
@@
Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IH}}({\bm{x}})\neq 0$. Note that $\frac{\partial\mathcal{L}_{\text{IH}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IH}}({\bm{x}})=0$. The detailed process for deriving this is provided in the Appendix A
@@
The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly demonstrates how the Inverted Hinge loss addresses the shortcomings of GA in knowledge unlearning for LLMs. First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with the opposite sign). This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly and intensively but also prevents spreading out of gradients for unlearning. During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$. Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease. This not only prevents unnecessary forgetting but also results in a bounded loss form.
3.4 FLoRA: Novel LoRA Initialization for LLM Unlearning
@@
Motivation. Aside from reduing the number of trainable parameters and thus reducing the memory requirements of LLM fine-tuning, LoRA serves as a potent form of regularization. While it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios (Biderman et al., 2024). Under the task of LLM unlearning, however, our objective strictly requires to completely remove knowledge of ${\mathcal{D}}_{f}$, and the strong stability from LoRA imposes a strong burden: vanilla LoRA unlearning requires large number of iterations through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in our Figure 2). Inspired by previous work on PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are important to ${\mathcal{D}}_{f}$ but not so important to ${\mathcal{D}}_{r}$ would benefit the unlearning process in terms of both convergence and post-unlearning performance. We hence design a novel LoRA initialization technique that measures the Fisher information on each weight ${\bm{W}}$ targeted for LoRA, and perform weighted low-rank decomposition from which we use the principal components to initialize adapter weights ${\bm{A}}$ and ${\bm{B}}$.
@@
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information dataset ${\mathcal{D}}$ provides on model parameters $\theta$. More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq. 3.4). However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead which, in the context of LLMs, can be computed as the average of squared gradients propagated from the language modeling objective (i.e. cross entropy loss):
@@
${\bm{F}}_{\theta}=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{\color[rgb]{0,0,0}{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta},$ (2)
@@
where $\mathcal{L}_{\text{LM}}$ is the next-token prediction loss of LM, such as $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$. Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that the next-token prediction loss on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
@@
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{rel}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to first identify parameters that are important specifically for ${\mathcal{D}}_{f}$, but unimportant to ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus choose to use the relative Fisher information matrix rather than $\hat{{\bm{F}}}^{f}$.
@@
Fisher-weighted LoRA Initialization (FLoRA). Intuitively, initializing LoRA weights to be parameters that are important in generating ${\mathcal{D}}_{f}$ would benefit unlearning as gradients would only modify parameters important to ${\mathcal{D}}_{f}$ while keeping all other parameters intact. Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{rel}$ for each target parameter ${\bm{W}}$, we propose to initialize LoRA adapter weights as the solution to the following problem of Weighted Low-Rank Approximation (WLRA):
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}$ (3)
@@
Note that when all weights equal one, this problem reduces down to a standard low-rank approximation problem, of which the solution can easily be computed via rank-$r$ SVD. In general, however, this minimization problem does not have a closed-form solution and thus requires iterative optimization (Srebro & Jaakkola, 2003). While we may resort to iterative methods towards initializing LoRA weights in a fine-grained manner, this would undermine the gains in efficiency due to employing LoRA. Thus, we tackle a simplified version of WLRA instead by assuming that parameters in each row of ${\bm{W}}$ share the same importance, for which we use the square-root of the row-wise sum of the relative Fisher information matrix as
@@
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\text{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\ \bm{\cdot}\ \mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$ (4)
@@
where $\mathds{1}\in\mathbb{R}^{k}$ is the all-one vector. $\text{diag}(\cdot)$ and $\bm{\cdot}$ denote the vector diagonalization function and the matrix-vector product, respectively. Unlike general WLRA, the row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}{\bm{1}})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
@@
Given the solution, we use the solution ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the output is not perturbed after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}{\bm{A}}$. In summary, our Fisher-Weighted initialization extracts parameters that are important only to ${\mathcal{D}}_{f}$ and not to ${\mathcal{D}}_{r}$ such that LoRA tuning can be focused towards erasing knowledge of ${\mathcal{D}}_{f}$.
3.5 Final Loss Function for Unlearning
Following initialization with FLoRA, unlearning is performed on the model $\Theta=\theta\cup\theta_{FLoRA}$, which consists of the original weights $\theta$ and the LoRA weights for each layer $\theta_{FLoRA}=\{{\bm{A}}^{*}_{l},{\bm{B}}^{*}_{l}\}_{l=1}^{L}$, where $L$ represents the number of layers where FLoRA is applied. During unlearning process, $\theta$ remains frozen, and only updates to the FLoRA weights $\theta_{FLoRA}$ are made. Furthermore, when performing unlearning on the forget dataset $D_{f}$, we incorporate the auxiliary retain set $D_{r}$, which contains general knowledge (e.g., the Wikitext dataset), as a default setting, following the gradient difference approach (Liu et al., 2022a; Maini et al., 2024). The final loss function using the proposed Inverted Hinge Loss is defined as follows:
$\operatorname*{minimize}_{\theta_{FLoRA}}\sum_{{\bm{x}}_{r}\in D_{f},{\bm{x}}_{f}\in D_{r}}\mathcal{L}_{\text{IH}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (5)

### v2

@@ 3.2 Preliminary Results
Despite its wide use in domain adaptation and instruction tuning, LoRA is not yet explored under the task of LLM unlearning to the best of our knowledge. Therefore, we first share empirical results from low-rank adapting LLMs using GD as our objective to motivate our approach. Figure 2 shows the results. Notably, vanilla LoRA suffers from lack of plasticity and ends up failing to sufficiently unlearn ${\mathcal{D}}_{f}$ within 20 epochs. When running more unlearning epochs or increasing the learning rate for sufficient unlearning, the model loses its previously acquired reasoning and generative capabilities, as shown in the significant decrease in Reasoning and Dialogue performances. In the remainder of this section, we present two techniques towards making LLM unlearning viable while enjoying the efficiency of LoRA.
3.3 Inverted Hinge Loss: A Novel Loss Function for LLM Unlearning
@@
Motivation. We analyze the inherent issues of GA from the perspective of its derivative. The output layer of a language model is a softmax layer that outputs probabilities over the vocabulary. Let ${y}_{t}$ be the logits (pre-softmax activations) produced by the LLM model for the $t$-th token, and let $V$ be the vocabulary size. The probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function: $p_{\theta}(x_{t}|x_{<t})={\exp(y_{t}^{(x_{t})})}/{\sum_{v=1}^{V}\exp(y_{t}^{(v)})}$ where $y_{t}^{(x_{t})}$is the logit corresponding to the true token $x_{t}$ and $y_{t}^{(v)}$ is the logit corresponding to the $v$-th token in the vocabulary. When we use $\mathcal{L}_{\text{GA}}$ for unlearning for LLMs, the gradient of the log-probability with respect to the logits is
@@
$\frac{\partial\log\left(p_{\theta}(x_{t}|x_{<t})\right)}{\partial y_{t}^{(v)}}=\begin{cases}1-p_{\theta}(x_{t}|x_{<t})&\text{if }v=x_{t}\\ -p_{\theta}(v|x_{<t})&\text{if }v\neq x_{t}\end{cases}$
From this derivative of GA, we can interpret its unlearning mechanism: GA decreases the prediction score corresponding to the true token $x_{t}$, while increasing the prediction score for all other tokens $v\neq x_{t}$ in the vocabulary. Consequently, we hypothesize that GA suffers from the following problems: 1) the objective is unbounded with the possibility of divergence during optimization, 2) gradient updates are inefficient due to gradients spreading across all other logits with large vocabulary sizes, and 3) there is unnecessary degradation of generative performance due to increasing the logits for all other tokens regardless of textual coherence.
@@
Inverted Hinge Loss. To cope with aforementioned limitations of GA, we aim to design a new loss function that achieves effective unlearning by decreasing the prediction score of the true token, while focusing gradient updates on only a minimal number of viable replacements for the ground-truth token. Inspired by the Hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge loss (IHL) as below:
@@
As the probability $p_{\theta}(x_{t}|x_{<t})$ is given by the softmax function, the derivative of $\mathcal{L}_{\text{IHL}}({\bm{x}})$ with respect to $y_{t}^{(v)}$ is:
@@
Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IHL}}({\bm{x}})\neq 0$. Note that $\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IHL}}({\bm{x}})=0$. The detailed derivation can be found in Appendix A.
@@
The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly illustrate how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs. First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign). This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly but also prevents spreading out of gradients. During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$. Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease. This not only prevents unnecessary forgetting but also results in a bounded loss function, allowing stable unlearning via stochastic gradient descent.
3.4 FILA: a Novel LoRA Initialization for LLM Unlearning
@@
Motivation. Biderman et al. (2024) discovered that, while it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios. While this stability is beneficial to retaining knowledge of ${\mathcal{D}}_{r}$ during unlearning, it also imposes a strong burden under the strict objective to remove knowledge of ${\mathcal{D}}_{f}$ completely. As a result, vanilla LoRA unlearning requires large number of model updates through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in Figure 2). Drawing inspiration from PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ would reinforce the model’s plasticity to forgetting ${\mathcal{D}}_{f}$, thereby accelerating the rate of unlearning and minimizing catastrophic forgetting on ${\mathcal{D}}_{r}$. In order to quantify the importance of parameters to particular data points, we leverage Fisher information matrices.
@@
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information a dataset ${\mathcal{D}}$ provides on model parameters $\theta$. More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq. 2). However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead. In the context of LLMs, this can be computed as:
@@
${\bm{F}}_{\theta}({\mathcal{D}})=\mathbb{E}_{\mathcal{D}}\left[\left(\dfrac{\partial}{\partial\theta}\log p{{}_{\theta}}({\mathcal{D}}|\theta)\right)^{2}\right]\approx\dfrac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}\in{\mathcal{D}}}\left(\dfrac{\partial}{\partial\theta}\mathcal{L}_{\text{LM}}({\bm{x}};\theta)\right)^{2}\eqcolon\hat{{\bm{F}}}_{\theta}({\mathcal{D}}),$ (2)
@@
where $\mathcal{L}_{\text{LM}}$ denotes the next-token prediction loss used to pretrain LMs, $\mathcal{L}_{\text{LM}}({\bm{x}};\theta)=\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$. Within our LLM unlearning setup, a high empirical Fisher information measured with ${\mathcal{D}}_{f}$ indicates that $\mathcal{L}_{\text{LM}}$ on ${\mathcal{D}}_{f}$ leads to large absolute gradients on the parameter under concern, and we consider such parameters to be important in generating sequences in ${\mathcal{D}}_{f}$.
@@
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}\coloneqq\hat{{\bm{F}}}_{\bm{W}}({\mathcal{D}}_{f})$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus use $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ rather than $\hat{{\bm{F}}}_{\bm{W}}^{f}$.
@@
Fisher-weighted Initialization of Low-rank Adapters. Given the relative importance $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ for each target weight ${\bm{W}}$, we propose to initialize the corresponding LoRA adapter weights with the solution to the following Weighted Low-Rank Approximation (WLRA) problem:
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\sum_{i,j}\left([\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}({\bm{W}}-{\bm{B}}{\bm{A}})_{i,j}\right)^{2}.$
@@
Note that when all weights $[\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}]_{i,j}$ equal one, WLRA reduces to standard low-rank matrix approximation, for which the solution can easily be computed via rank-$r$ SVD. For general weights, however, this minimization problem does not have a closed-form solution and requires iterative optimization (Srebro & Jaakkola, 2003). While we may resort to iterative methods to initialize LoRA weights in a fine-grained manner, this would undermine the efficiency gains from deploying low-rank adapters. Therefore, we assume that parameters in each row of ${\bm{W}}$ share the same importance equal to the square-root of the row-wise sum of $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$, and simplify the problem to
@@
$\min_{{\bm{A}}\in\mathbb{R}^{r\times k},{\bm{B}}\in\mathbb{R}^{d\times r}}\left\|\texttt{diag}\left((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{\frac{1}{2}}\right)({\bm{W}}-{\bm{B}}{\bm{A}})\right\|_{2}$
@@
with $\mathds{1}\in\mathbb{R}^{k}$ and $\texttt{diag}(\cdot)$ indicating the all-one vector and the vector diagonalization function, respectively. Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
@@
Given this solution, we use ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the model behavior remains the same after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}^{*}{\bm{A}}^{*}$. Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$, but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
3.5 Final Loss Function for LLM Unlearning
In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$, consisted of original pretrained weights $\bm{\theta}$ and the FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA. Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$. The final loss function using both the proposed IHL and FILA is defined as follows:
$\operatorname*{minimize}_{\theta_{\text{FILA}}}\sum_{{\bm{x}}_{r}\in{\mathcal{D}}_{f},{\bm{x}}_{f}\in{\mathcal{D}}_{r}}\mathcal{L}_{\text{IHL}}({\bm{x}}_{f})+\mathcal{L}_{\text{LM}}({\bm{x}}_{r})$ (3)